OpenStudy (anonymous):

Nathan stores coffee powder in a steel container with a capacity of 68.75 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 3.5 inches, what is its radius? Answer 2.5 inches 3.5 inches 5.0 inches 7.0 inches

OpenStudy (anonymous):

OK, so you're given volume V=68.75 and height h=3.5 and you want to find radius r. Do you remember the formula for the volume of a cylinder?

OpenStudy (anonymous):

pi r^2 h

OpenStudy (anonymous):

right, V=pi*r^2*h see any letters you can substitute with the given numbers?

OpenStudy (anonymous):

You can plug in numbers for V and h. After doing so, you can solve for r. If you need more help, ask.

OpenStudy (anonymous):

zax97, do you know the definition of congruent?

OpenStudy (anonymous):

\[\pi r^2 h=\frac{275}{4} \]The expression above is the steel cylinder's volume. Replace h with the value of 7/2 and solve for r. The exact symbolic answer is\[5 \sqrt{\frac{11}{14 \pi }} \]The decimal numbers have been changed to their fractional equivalents. Use a calculator to convert the symbolic answer to a decimal number.

OpenStudy (anonymous):

congruent: "coinciding exactly when superimposed."